OpenStudy (anonymous):

what is b^3 = 1/64, find b

Directrix (directrix):

b^3 = 1/64 b^3 = (1/4)^3 Can you take it from there?

OpenStudy (campbell_st):

take the cube root of both sides of the equation

OpenStudy (anonymous):

but how?

OpenStudy (anonymous):